OpenStudy (solomonzelman):

remind me please, I indeed feel stupid to ask this, but would the 3x^2' be 6x ?

OpenStudy (nincompoop):

power rule

OpenStudy (austinl):

It would indeed. \(\large{\frac{d}{dx}ax^n=a\cdot nx^{n-1}}\)

OpenStudy (nincompoop):

X^n = nX^(n-1)

OpenStudy (solomonzelman):

Sorry for editing too much ((

OpenStudy (nincompoop):

there's a prime after 2?

OpenStudy (turingtest):

As a footnote, the correct English phrasing would be"how do I take the derivative of...?". To derive something has a different meaning.

OpenStudy (turingtest):

I don't think that deserved two medals, but ok :P

OpenStudy (solomonzelman):

Yeah, ik, also correct way would be to capitalize word how, but anyway, you know what I mean, and I can tell....

OpenStudy (nincompoop):

f(x) = 3x^2; f' 2x*dx

OpenStudy (solomonzelman):

So would I just say \[\huge\color{blue}{ ax^n=a \times n x ^ {( n -1) } } \] (where a*n is not powered to n-1 )

OpenStudy (turingtest):

you definitely don't want to say that, since it's not true

OpenStudy (austinl):

That is an equation, the \(derivative\) of the left side is the right side.

OpenStudy (turingtest):

\[\huge\color{blue}{\frac d{dx} ax^n=a \times n x ^ {( n -1) } }\]that d/dx out front makes all the difference, and is *not* to be trivially ignored, as perhaps the wording is.

OpenStudy (nincompoop):

true you have to denote what you are doing, people use d/dx and some use primes. use whichever suits you

OpenStudy (solomonzelman):

is d/dx just saying that you are deriving?

OpenStudy (nincompoop):

the d denotes what you are doing, the dx at the numerator tells you with respect to what, in your case the x

OpenStudy (austinl):

For example, if you have a function, \(y=5x+7\) \(\large{\frac{dy}{dx}}\) means that you are taking the derivative of the function y with respect to x.

OpenStudy (nincompoop):

learn calculus without the jargons http://finedrafts.com/files/CUNY/math/calculus/S%20Thompson/

OpenStudy (turingtest):

if we have \[y=f(x)\] then the derivative of the function with respect to x is denoted by one of the following\[\frac{dy}{dx}=y'=f'(x)=D_x f(x)\]

OpenStudy (nincompoop):

it does not make use of limits and it shows you what differentiation really is like.

OpenStudy (solomonzelman):

Alright, am in precalc and not up to derivatives, although I know L'H'S and little more staff. Weird ik :) But how would I put in a second derivative, like using walpharampha (or whatever it is, the online calc) would I say dy^2/dx ?